What maximum height attained by the object that is projected vertically upward with a
speed of 60 m/s? How long will it take for the object to reach the maximum height?

3. A ball is batted into the air and the ball found at a point 200 m distance horizontally in
5.0 seconds. Neglecting air resistance. Find the maximum height attain by the ball.
4. A Cannon is fired with a muzzle velocity of 300 m/s, at an angle of 60 degrees. Find the
range, time of flight, maximum height attained by the projectile.
1
Expert's answer
2021-10-11T08:56:02-0400

The maximum height and time is

"H=\frac{v^2}{2g}=184\text{ m}.\\\space\\
t=\frac{v}{g}=6.12\text{ s}."

3. The height is

"H=\frac{g(t/2)^2}{2}=31\text{ m}."

4. Use similar equations:


"H=\frac{(v^2\sin^2\theta)}{2g}=3444\text{ m}.\\\space\\
T=2\frac{v\sin\theta}{g}=53\text{ s}.\\\space\\
R=Tv\cos\theta=7950\text{ m}."